OpenStudy (anonymous):

Factorize : \[x^4+ 3x^2+ 2\]

OpenStudy (anonymous):

@Hero

Parth (parthkohli):

Take \(u = x^2\).\[u^2 + 3u +2\]

OpenStudy (ajprincess):

Assume \(x^2=y\) Then u will have \(y^2+3y+2\) Can u factorise this? @Glen_McGrath

Parth (parthkohli):

Or you could just do\[x^4 + 3x^2 + 2 = x^4 + x^2 + 2x^2 + 2 = x^2(x^2 + 1) + 2(x^2 + 1) = (x^2 + 2)(x^2 +1)\]

Parth (parthkohli):

Do you get it? @Glen_McGrath

OpenStudy (anonymous):

GOOD parth It matches my answer key

Parth (parthkohli):

yay!

OpenStudy (anonymous):

thanks again

Parth (parthkohli):

No problem. :-)